# Unleashing HyDRa: Hybrid Fusion, Depth Consistency and Radar for Unified 3D Perception

### v2

@@ Abstract
Low-cost, vision-centric 3D perception systems for autonomous driving have made significant progress in recent years, narrowing the gap to expensive LiDAR-based methods. The primary challenge in becoming a fully reliable alternative lies in robust depth prediction capabilities, as camera-based systems struggle with long detection ranges and adverse lighting and weather conditions. In this work, we introduce HyDRa , a novel camera-radar fusion architecture for diverse 3D perception tasks. Building upon the principles of dense BEV (Bird’s Eye View)-based architectures, HyDRa  introduces a hybrid fusion approach to combine the strengths of complementary camera and radar features in two distinct representation spaces. Our Height Association Transformer module leverages radar features already in the perspective view to produce more robust and accurate depth predictions. In the BEV, we refine the initial sparse representation by a Radar-weighted Depth Consistency. HyDRa  achieves a new state-of-the-art for camera-radar fusion of 64.2 NDS ($+1.8$) and 58.4 AMOTA (+1.5) on the public nuScenes dataset. Moreover, our new semantically rich and spatially accurate BEV features can be directly converted into a powerful occupancy representation, beating all previous camera-based methods on the Occ3D benchmark by an impressive 3.7 mIoU. Code and models are available at https://github.com/phi-wol/hydra.
1 Introduction
@@
Reliable and affordable 3D perception is an important cornerstone for safe and efficient operation of autonomous vehicles, enabling agents to navigate dynamic environments and complex scenarios. In this rapidly evolving field, deep learning-based Camera-LiDAR fusion has emerged as the de-facto standard for 3D reconstruction from multiple sensors, with a proven track record in 3D object detection [77, 45, 37], semantic segmentation [45, 12, 49], and object tracking [8, 37].
@@
While delivering highly accurate geometry sensing, LiDAR sensors still pose a high-cost barrier for large-scale deployment. In order to democratize access to autonomous driving systems, LiDAR-free, and vision-centric systems have gained significant traction in recent years and months. Especially with the advent of Bird’s Eye View (BEV) based architectures [56], camera-only methods have achieved remarkable progress [33, 67, 62]. Cameras provide dense and high-resolution semantic information, theoretically enabling the detection of objects at far distances. However, the main practical challenge lies in preserving the spatial information after 2D projection, leading to unstable depth estimation and localization errors [33, 55]. Another difficulty of passive sensors is posed by dynamic objects that are occluded, yet remain relevant for safe decision making.
@@
Still underrepresented in this research domain is the incorporation of radar sensors. Especially camera-radar fusion has been neglected by the academic community [12]. As low-frequency active sensors based on time-of-flight principles, they exhibit a remarkable resilience to adverse weather and lighting conditions, offering metric measurements for perception ranges up to 300 meters [78]. Furthermore, by leveraging the Doppler effect, radar reflections can yield valuable velocity information about moving and occluded objects, underscoring the potential benefits as a camera-complementary sensor. Combining the two sensor types is key to unlocking the full potential of accessible 3D perception systems [60].
Nevertheless, the task of integrating both sensor types into a coherent and robust perception system is a challenging, non-trivial task. Effectively associating radar detections with camera features is complicated by the unique characteristics of the radar data. The absence of elevation information, the sparse nature of the radar returns, the presence of noise from multi-path reflections, the occurrence of false positives, and the overall high measurement uncertainty make radar data difficult to process and associate with camera features [23, 60]. Consequently, the direct application of state-of-the-art LiDAR-like architectures on inherently sparse radar data has proven to be ineffective [51]. Despite being in point-cloud form, Yan et al. highlight that CMT [74], among the current leading methods on the nuScenes 3D object detection task, degrades in performance when incorporating the radar data into their pipeline. For this reason, we see the need for a significant shift in fusion paradigms when adopting radar. Traditionally, fusion approaches have been heavily biased towards LiDAR, with the bulk of the network’s capacity devoted to processing LiDAR point clouds, and camera data merely serving as a supplementary source of semantic information [65, 31, 19]. Leaving the LiDAR sensor out of the equation, new approaches significantly prioritize extracting dense semantic information from camera feeds and its projection into the BEV grid. The methodology relies heavily on advanced 2D feature extraction techniques [26, 43, 44] along with efficient view transformation and lifting modules [17, 34, 36]. In this new configuration, radar data assumes a role akin to the previous camera-based methods, enhancing the primary modality with sparse metric information and becoming a guidance module [12, 23].
Modern practices have focused on adaptive fusion in the BEV, reducing spatial misalignment during feature aggregation [45, 23]. Despite these improvements, they still struggle with robust depth estimation and rely partly on monocular depth cues for feature transformation [33]. Existing techniques, while innovative, do not fully capitalize on radar’s potential for unified robust depth sensing, often resulting in a dual-stage projection process where each modality stage relies on its own, potentially inconsistent, depth estimation. We argue that to fully unlock the potential, we must move the fusion stage even earlier. In this work, we propose HyDRa, a state-of-the-art camera-radar fusion architecture well equipped to tackle a variety of 3D perception tasks. Our contributions and important stepstones to enable LiDAR-free autonomous vehicles and advanced driver assistance systems are the following:
@@
Figure 1: Overall Architecture of HyDRa: Modality-specific backbones extract multi-view image and radar pillar features. The features are fused in two representation spaces: Perspective View and BEV-Space. 1. The radar features are associated with the image features by the Height Association Transformer. With the resulting radar-informed dense depth distribution, the forward projection module generates a sparse BEV representation. 2: The splatted semantic BEV features and radar-BEV features are concatenated and fused by a SE block. 3: A depth-aware backward projection refines this representation, guided by radar attention weights before being distributed to task-specific heads.
@@
    We introduce our Height Association Transformer (HAT) module and address the limitations of previous BEV-generating depth networks. By creating unified geometry-aware features in the perspective view, we significantly reduce the translation error of existing methods.
@@
    We upgrade the concept backprojection-as-refinement and establish a Radar-weighted Depth Consistency (RDC) for enhancing sparse fusion features in the BEV, tackling the issue of misaligned or unassociated features and occluded objects.
@@
    We extensively ablate our design choices and show that each of our contributions significantly improves the performance of our model on the nuScenes detection and tracking challenges [2], but the synergy of all components is crucial for achieving best-in-class results.
@@
2 Related Work
2.1 Camera-based Architectures
Dense BEV-based Architectures. Shifting from perspective-view 3D object detection [50, 86, 59, 68, 69] to Birds-Eye-View (BEV) based perception [56, 57], the BEV-based architectures have become the new standard for 3D object detection and tracking [28]. With the introduction of differentiable lifting methods, the image-based output of a standard 2D convolutional neural networks [13, 44] or a transformer-based feature extractor [9, 43] is transformed into the more downstream task-friendly BEV ground plane representation [86, 15]. The BEVDet series [18, 16, 17] and following work [45, 33] have raised the bar for efficient and powerful forward projection pipeline, predicting an explicit, dense, and pixel-wise depth distribution and pooling the camera-frustum features in the BEV space. BEVDepth [33], SoloFusion [55], and BEVStereo [32] focused on localization by leveraging LiDAR supervision during training, investigating long-term 4D fusion and temporal stereo matching for better depth estimation. BEVFormer [34, 75] has shown that inverse modeling of the view transformation[58] is a valid alternative to the forward projection pipeline; by back-projecting a set of pre-defined BEV grid queries. Spatial deformable cross attention [64, 88] implicitly learns to model the 3D relations by interacting with regions of interest across multiple image feature maps.
FB-BEV [36] combines both approaches into the forward-backward projection paradigm, modeling a more precise projection relationship. It addresses the sparsity and inaccuracy inherent in both projection types by ensuring a Depth Consistency between the two modules. We will build upon this promising concept and extend it to the multi-modal domain by introducing a novel Radar-weighted Depth Consistency and incorporating radar in complementary representation spaces to enforce a more accurate depth estimation.
Sparse BEV-based Architectures. Instead of modeling the complete surrounding, a concurrent line of research has established the modeling of the scene as a sparse set of object queries [4, 27, 82], decoding 3D proposals via implicit attention [41, 70, 42, 66]. In addition to long-term temporal fusion [55], StreamPETR introduces an object-level temporal mechanism and query denoising [27, 82] to establish an efficient and powerful detection and tracking paradigm [67]. SparseBEV reformulates the design by scale-adaptive self-attention and spatio-temporal sampling inspired by AdaMixer[11]. Without explicit dense BEV feature construction, the model excels in detecting foreground objects, with comparable higher translation errors [55]. These sparse architectures cannot be easily applied for full environmental understanding, such as the drivable area and road boundaries [62, 61]. Hence, we focus on dense BEV-based architectures, as they are best suited for the full stack of planning-oriented end-to-end autonomous driving.
2.2 Multi-modal Architectures
@@
Initial works on camera-radar fusion have been restricted by the view disparity between the sensor inputs. GrifNet [21], CRAFT [22] and Centerfusion [51] typically follow Region-of-Interest-pooled late-fusion approaches in the perspective or BEV-view. They focus on producing camera-based detections, refined by the offsets to associated radar point detections. Others directly concatenate view-aligned radar feature maps to high-level image features and feed the fused features into a standard 3D detector head [72, 46]. They rely on simple heuristics-based radar point duplication due to the missing height information. We address these shortcomings by designing an end-to-end trainable association to image feature height dimension. X3KD [24] investigates a knowledge distillation framework distilling across modalities from a LiDAR-based teacher, whereas BEVGuide [49] and FUTR3D [5] employ BEV query-based feature sampler with a transformer decoder, but none of them leverages the radar at all for view transformation.
In general, fusing the radar in its natural 3D representation form of the BEV grid has proven to be most effective sofar [78, 12, 23]. RCBEV [85] empirically showed that using specialized or heavy point-cloud-processing feature backbones does not bring performance gains - hence why our state-of-the-artarchitecture leverages efficient pillar-encoding voxelization [25, 29].
The former state-of-the-art in radar-camera fusion, CRN [23] employs the same underlying architecture as BEVDepth [33]. Kim et al. introduce a multi-modal deformable attention mechanism, a BEV-fusion feature operator with a bigger receptive field than simple concatenation [45]. Additionally, they propose a radar-assisted view transformation (RVT) to complement a monocular depth estimation network with a radar-occupancy-based view projection, working in two separate streams. They were the first on nuScenes [2] to show that with an increasing perception range of up to 100m, radar-fusion-based methods outperform LiDAR at far distances [80]. While CRN’s RVT is a step in the right direction, it still relies on two parallel but separate view transformations. We propose a more unified approach to solve this spatial misalignment introduced by monocular depth cues by moving the fusion to a lower level and designing a hybrid fusion.
2.3 Occupancy Prediction
Occupancy Prediction is a new task emerging in the context of 3D perception with the goal of predicting the occupancy and semantic class of every voxel in the surrounding 3D environment [62, 61] It is an important intermediate task in the full stack of planning-oriented end-to-end autonomous driving, positively affecting planning L2 error and collision rate. Semantic occupancy prediction even helps to outperform LiDAR-based planners [15]. Dense architectures (2.1) are well equipped to predict the 3D status of dynamic and static foreground as well as background voxels, as this task is heavily dependent on the quality of depth estimation [35, 53, 81, 48]. Instead of working on the 2D BEV space, the ground plane grid extends to 3D voxels, which are then predicted by 3D convolutional encoders. We show, that a camera-radar fused BEV representation is the best suited for this task, excelling over all previous camera-only methods.
3 HyDRa Architecture
@@
  • BEV Fusion: The forward projection module [18] generates the initial BEV representation and is concatenated with the radar pillars channels and fused by a simple Squeeze-and-Exictation layer [14, 37]. Cached history features [55] can be optionally fused.
@@
BEVDepth [33] C R101 $512\times 1408$ 53.5 41.2 0.565 0.266 0.358 0.331 0.190
BEVGuide [49] C+R R101 $224\times 480$ 53.7 42.1 - - - 0.390 -
SparseBEV [39] C R101 $512\times 1408$ 59.2 50.1 0.562 0.265 0.321 0.243 0.195
StreamPETR [67] C R101 $512\times 1408$ 59.2 50.4 0.569 0.262 0.315 0.257 0.199
@@
HyDRa C+R R101 $512\times 1408$ 61.7 53.6 0.416 0.264 0.407 0.231 0.186
Table 1: 3D Object Detection on nuScenes val set. ‘L’, ‘C’, and ‘R’ represent LiDAR, Camera, and Radar, respectively. ( Backbones benefitted from perspective depth pre-training.)
3.1 Vision-Centric Foundation
In our method we propose an extension to existing vision-centric 3D perception systems. Thus we build upon the recent advancements in camera-based high-performance object detection (BEVDet Series [18, 33]) and the simple but effective concept of fusing radar and image features in the BEV space [12]. The setup follows our baseline method CRN [23]. The image encoder, multi-resolution feature pyramid neck [38], the Context-Net, Depth-Net, and BEV-Encoder [13] follow the same architecture choices as previous work [18, 33].
3.2 Height Association Transformer
Our method utilizes the strength of each modality to overcome the main challenge in vision-centric depth prediction. We propose a new transformer-based plug-in module to leverage the complementary radar pillar features (cf. Fig. 3). The primary challenge is associating the radar features $R$ with the respective image features $F$ in the perspective view. For comparison, CRN’s RVT [23] collapses the image features due to missing information in the height dimension and uses the sparse encoded radar-only occupancy in the frustum to splat the features in the BEV space. Instead of using a separate view transformation or the noisy depth map as input [46], we aim to learn which part of the downsampled feature map should benefit from the noisy and ambiguous radar encoding with an extended receptive field along the complete image height.
Let $F\in\mathbb{R}^{B\times N\times H\times W\times C}$ denote 2D image backbone output, while $R\in\mathbb{R}^{B\times N\times 1\times W\times C}$ describes the radar pillar features in the respective camera frustum. $B$ is the batch size, $N$ the number of camera views, $H$ the downsampled feature height dimension, $W$ the downsampled feature width, $D$ the number of depth bins in the camera frustum and $C$ the same-size embedding dimension. For efficiency, we query a single column $w$ of discrete monocular height features $F_{bnw}\in\mathbb{R}^{H\times 1\times C}$ for the $w$-th respective radar feature keys and values $R_{bnw}\in\mathbb{R}^{1\times D\times C}$ on the ground plane. Therefore, we reshape the feature tensors to a new batch size $B^{\prime}=B\times N\times W$. For each ’sequence’ we apply a learnable positional embedding. Next, we encode the height bins with self-attention and decode and fuse the radar features via cross-attention [64]. A peak in activations on the depth bin plane is compared with promising height proposals, encoding the geometric relationship in the attention, inducing sparse but strong metric cues into the dense feature space. Either modality is missing crucial information that the other sensor can provide. To stabilize the learning, we apply a zero-initialized layer norm [1] to the fusion features, and add them residually to the original image features. The resulting feature map $F^{\prime}$ is the input for the depth and context network, producing the final more robust depth distribution.
@@ 3.2 Height Association Transformer
PointPillars [25] L Pillars 55.0 40.1 0.392 0.269 0.476 0.270 0.102
CenterPoint [80] L Voxel 67.3 60.3 0.262 0.239 0.361 0.288 0.136
KPConvPillars [63] R Pillars 13.9 4.9 0.823 0.428 0.607 2.081 1.000
CenterFusion [51] C+R DLA34 44.9 32.6 0.631 0.261 0.516 0.614 0.115
MVFusion [72] C+R V2-99 51.7 45.3 0.569 0.269 0.410 0.781 0.203
CRAFT [22] C+R DLA34 52.3 41.1 0.467 0.268 0.456 0.519 0.114
BEVDepth [33] C ConvNeXt-B 60.9 52.0 0.445 0.243 0.352 0.347 0.127
SOLOFusion [55] C ConvNeXt-B 61.9 54.0 0.453 0.257 0.376 0.276 0.148
BEVFormerV2 [75] C InternImage-B 62.0 54.0 0.47.9 0.248 0.320 0.295 0.123
CRN [23] C+R ConvNeXt-B 62.4 57.5 0.416 0.264 0.456 0.365 0.130
StreamPETR [67] C V2-99 63.6 55.0 0.479 0.239 0.317 0.241 0.119
SparseBEV [39] C V2-99 63.6 55.6 0.485 0.244 0.332 0.246 0.117
HyDRa C+R V2-99 64.2 57.4 0.398 0.251 0.423 0.249 0.122

### v3

# for Unified 3D Perception

@@ Abstract
Low-cost, vision-centric 3D perception systems for autonomous driving have made significant progress in recent years, narrowing the gap to expensive LiDAR-based methods. The primary challenge in becoming a fully reliable alternative lies in robust depth prediction capabilities, as camera-based systems struggle with long detection ranges and adverse lighting and weather conditions. In this work, we introduce HyDRa, a novel camera-radar fusion architecture for diverse 3D perception tasks. Building upon the principles of dense Bird’s-Eye-View (BEV)-based architectures, HyDRa introduces a hybrid fusion approach to combine the strengths of complementary camera and radar features in two distinct representation spaces. Our Height Association Transformer module leverages radar features already in the perspective view to produce more robust and accurate depth predictions. In the BEV, we refine the initial sparse representation by a Radar-weighted Depth Consistency. HyDRa achieves a new state-of-the-art for camera-radar fusion of 64.2 NDS (+1.8) and 58.4 AMOTA (+1.5) on the public nuScenes dataset. Moreover, our new semantically rich and spatially accurate BEV features can be directly converted into a powerful occupancy representation, beating all previous camera-based methods on the Occ3D benchmark by an impressive 3.7 mIoU. Code and models are available at https://github.com/phi-wol/hydra .
I Introduction
@@
Reliable and affordable 3D perception is an important cornerstone for safe and efficient operation of autonomous vehicles, enabling agents to navigate dynamic environments and complex scenarios. In this rapidly evolving field, deep learning-based Camera-LiDAR fusion has emerged as the de-facto standard for 3D reconstruction from multiple sensors [1, 2, 3, 4, 5].
@@
While delivering highly accurate geometry sensing, LiDAR sensors still pose a high-cost barrier for large-scale deployment. In order to democratize access to autonomous driving systems, LiDAR-free, and vision-centric systems have gained significant traction in recent years and months [6, 7, 8]. However, the main practical challenge lies in preserving the spatial information after 2D projection, leading to unstable depth estimation and localization errors [6, 9]. Another difficulty of passive sensors is posed by dynamic objects that are occluded, yet remain relevant for safe decision making.
@@
Still underrepresented in this research domain is the incorporation of radar sensors. Especially camera-radar fusion has been neglected by the academic community [10]. As low-frequency active sensors based on time-of-flight principles, they exhibit a remarkable resilience to adverse weather and lighting conditions, offering metric measurements for perception ranges up to 300 meters [11]. Furthermore, by leveraging the Doppler effect, radar reflections can yield valuable velocity information about moving and occluded objects. Combining the two sensor types is key to unlocking the full potential of accessible 3D perception systems [12].
Figure 1: Bridging the view disparity in BEVFusion [4], CRN [13], and our HyDRa. We leverage multi-modal feature fusion alraedy for depth splatting.
@@
Figure 2: Architecture of HyDRa: The modality-specific features are fused in two representation spaces: Perspective View and BEV-Space. 1. The radar features are associated with the image features by the Height Association Transformer. With the resulting radar-informed dense depth, the forward projection module generates a sparse BEV representation. 2. The splatted semantic BEV features and radar-BEV features are concatenated and fused.  3. A depth-aware backward projection refines this representation, guided by radar attention weights before being distributed to task-specific heads.
Nevertheless, effectively associating radar detections with camera features is complicated by the unique characteristics of the radar data. The absence of elevation information, the sparse nature of the radar returns, the presence of noise from multi-path reflections, and the overall high measurement uncertainty make radar data difficult to process and associate with camera features [13, 12]. Consequently, the direct application of state-of-the-art LiDAR-like architectures has proven to be ineffective [14]. CMT [15], among the current leading methods on the nuScenes benchmark, degrades in performance when incorporating radar points into their model. For this reason, we see the need for a significant shift in fusion paradigms when adopting radar.
Modern practices have focused on adaptive fusion in the BEV, reducing spatial misalignment during feature aggregation [2, 13]. Despite these improvements, they still struggle with robust depth estimation and rely partly on monocular depth cues for feature transformation [6] (cf. Fig. 1). Existing techniques do not fully capitalize on radar’s potential for unified robust depth sensing, often resulting in a dual-stage projection process where each modality stage relies on its own, potentially inconsistent, depth estimation. We argue that to fully unlock the potential, we must move the fusion stage even earlier.
In this work, we propose HyDRa, a state-of-the-art camera-radar fusion architecture well equipped to tackle a variety of 3D perception tasks. Our contributions are the following:
@@
    We introduce our Height Association Transformer (HAT) module and address the limitations of previous BEV-generating depth networks. By creating unified geometry-aware features in the perspective view, we significantly reduce the translation error.
@@
    We establish a Radar-weighted Depth Consistency (RDC) for enhancing sparse features in the BEV, tackling misaligned features and occluded objects.
@@
    We extensively ablate our design choices on the nuScenes and View-of-Delft detection and tracking challenges and show that the synergy of all components is crucial for achieving best-in-class results.
@@
II Related Work
II-A Camera-based Architectures
Shifting from perspective-view 3D object detection [17, 18, 19] to Birds-Eye-View (BEV) based perception [20, 21], the BEV-based architectures have become the new standard for 3D object detection and tracking [22, 23]. With the introduction of differentiable lifting methods, the image-based output of a standard 2D feature extractor [24, 25, 26, 27] is transformed into the more downstream task-friendly BEV ground plane representation [28, 29, 30]. The BEVDet series [31, 32, 33] and following works [2, 6, 9] have raised the bar for an efficient and powerful forward projection pipeline, predicting an explicit, dense, and pixel-wise depth distribution and pooling the camera-frustum features in the BEV space, while BEVFormer [34] introduces a spatial cross attention [35, 36] to implicitly model the 3D relations of regions of interest. FB-BEV [37] combines both approaches, addressing the sparsity and inaccuracy inherent in both projection types by ensuring a Depth Consistency between the two modules.
We will build upon this concept and extend it to the multi-modal domain by introducing a novel Radar-weighted Depth Consistency and incorporating radar in complementary representation spaces to enforce a more accurate depth estimation.
II-B Multi-modal Architectures
@@
Initial works on camera-radar fusion typically follow Region-of-Interest-pooled late-fusion approaches in the perspective or BEV-view [38, 14, 39]. Others directly concatenate view-aligned radar feature maps to high-level image features and feed the fused features into a standard 3D detector head [40, 41]. X3KD [42], BEVGuide [3] and FUTR3D [43] employ BEV query-based feature sampler with a transformer decoder. In general, fusing the radar in its natural 3D representation of the BEV grid has proven to be most effective sofar [11, 10, 44, 13]. RCBEV [44] empirically shows that using specialized or heavy point-cloud-processing feature backbones does not result in performance gains - hence why our state-of-the-art architecture leverages efficient pillar voxelization [45, 46].
CRN [13] introduces a multi-modal deformable attention mechanism, a BEV-fusion feature operator with a larger receptive field than simple concatenation [2]. While they rely on a monocular depth network and flatten image features to use the radar occupancy as a second view-transformation, we aim to strengthen the depth estimation itself by early-fusion of the two independent feature spaces. Our approach addresses spatial misalignment from monocular depth cues by moving fusion to a lower level, employing a hybrid fusion.
III HyDRa Architecture
@@
  • BEV Fusion: The forward projection module [31] generates the initial BEV representation and is concatenated with radar pillar channels and fused by a simple Squeeze-and-Exictation layer [47, 4].
@@
TABLE I: 3D Object Detection on nuScenes val set. ‘L’, ‘C’, and ‘R’ represent LiDAR, Camera, and Radar, respectively.
@@
III-A Vision-Centric Foundation
In our method we propose an extension to existing vision-centric 3D perception systems. Thus we build upon the recent advancements in camera-based high-performance object detection (BEVDet Series [31, 6]) and the simple but effective concept of fusing radar and image features in the BEV space [10]. The image encoder with a feature pyramid neck [50], the Context-Net, Depth-Net, and BEV-Encoder [24] follow the same architecture choices as previous work [31, 6].
III-B Height Association Transformer
Our method utilizes the strength of each modality to overcome the main challenge in vision-centric depth prediction. We propose a new transformer-based plug-in module to leverage the complementary radar pillar features (cf. Fig. 4). The primary challenge is associating the radar features $R$ with the respective image features $F$ in the perspective view. We aim to learn which part of the downsampled feature map should benefit from the ambiguous radar encoding with an extended receptive field along the complete image height (in principle reversing the view transformation).
Let $F\in\mathbb{R}^{B\times N\times H\times W\times C}$ denote 2D image backbone output, while $R\in\mathbb{R}^{B\times N\times 1\times W\times C}$ describes the radar pillar features in the respective camera frustum. $B$ is the batch size, $N$ the number of cameras, $H$ and $W$the feature height and width, $D$ the number of depth bins in the camera frustum and $C$ the same-size embedding dimension. For efficiency, we query a single column $w$ of discrete monocular height features $F_{bnw}\in\mathbb{R}^{H\times 1\times C}$ for the $w$-th respective radar feature keys and values $R_{bnw}\in\mathbb{R}^{1\times D\times C}$ on the ground plane. Therefore, we reshape the feature tensors to a new batch size $B^{\prime}=B\times N\times W$. For each ’sequence’ we apply a learnable positional embedding. Next, we encode the height bins with self-attention and decode and fuse the radar features via cross-attention [35]. A peak in activations on the depth bin plane is compared with promising height proposals, encoding the geometric relationship in the attention, inducing sparse but strong metric cues into the dense feature space. Either modality misses crucial information that the other sensor can provide. The resulting feature map $F^{\prime}$ is the input for the depth and context network, producing the final more robust depth distribution.